### 08 Jan 2024 v1

# An exploratory study on automatic identification of assumptions in the development of deep learning frameworks

Abstract
Stakeholders constantly make assumptions in the development of deep learning (DL) frameworks. These assumptions are related to various types of software artifacts (e.g., requirements, design decisions, and technical debt) and can turn out to be invalid, leading to system failures. Existing approaches and tools for assumption management usually depend on manual identification of assumptions. However, assumptions are scattered in various sources (e.g., code comments, commits, pull requests, and issues) of DL framework development, and manually identifying assumptions has high costs (e.g., time and resources). To overcome the issues of manually identifying assumptions in DL framework development, we constructed a new and largest dataset (i.e., AssuEval) of assumptions collected from the TensorFlow and Keras repositories on GitHub; explored the performance of seven traditional machine learning models (e.g., Support Vector Machine, Classification and Regression Trees), a popular DL model (i.e., ALBERT), and a large language model (i.e., ChatGPT) of identifying assumptions on the AssuEval dataset. The experiment results show that: ALBERT achieves the best performance (f1-score: 0.9584) of identifying assumptions on the AssuEval dataset, which is much better than the other models (the 2nd best f1-score is 0.6211, achieved by ChatGPT). Though ChatGPT is the most popular large language model, we do not recommend using it to identify assumptions in DL framework development because of its low performance on the task. Fine-tuning ChatGPT specifically for assumption identification could improve the performance. This study provides researchers with the largest dataset of assumptions for further research (e.g., assumption classification, evaluation, and reasoning) and helps practitioners better understand assumptions and how to manage them in their projects.
keywords:
Assumption, Automatic Identification, Deep Learning Framework, TensorFlow, Keras
journal: Elsevier
1 Introduction
Deep learning (DL) is a branch of machine learning (ML), which has been extensively used in various areas as a key component, such as image classification, question-and-answer, and speech recognition. To improve the efficiency of developing DL-based applications, many popular DL frameworks (e.g., Theano theano2022 , Caffe jia2014caffe , TensorFlow Abadi2016 , and PyTorch Steiner2019 ) have been designed, implemented, and used in DL systems. Due to lack of knowledge or information, time pressure, complex context, etc., various uncertainties emerge during the development of DL frameworks, leading to assumptions made in projects Yangsms2018 . The assumptions in DL frameworks are different from those of traditional software systems. In addition to assumptions about system requirements, design, and components, many assumptions in DL frameworks are about data, DL models, and training siebert2021 . For example, in the documents (e.g., multigpu.md) of Caffe, developers explicitly mention assumptions, such as assuming that GPUs for training are of the same type [^1] https://github.com/BVLC/caffe/blob/master/docs/multigpu.md. Erickson et al. pointed out that the assumptions of the input data dimensions in TensorFlow and Theano are different, and this should be considered in system design and coding when using such DL frameworks Erickson2017 .
The importance of assumptions and their management in software development (including DL framework development) has been highlighted in many studies. For example, Corbató Corbato1991 mentioned in his ACM Turing Award lecture that “design bugs are often subtle and occur by evolution with early assumptions being forgotten as new features or uses are added to systems.” Garlan et al. pointed out that incompatible assumptions in software architecture can cause architectural mismatch Garlan2009 . Lewis et al. also got similar results in ML systems: since there are different types of stakeholders (e.g., data scientist, software engineer, and system user) of an ML system, they could make different but incompatible or invalid assumptions, leading to system misunderstanding, mismatch, etc lewis2021 . Zhang et al. Zhang2020 collected 4,960 real failures of several DL frameworks, such as TensorFlow, PyTorch, and Keras, and found that part of system failures are caused by assumptions. For example, during the evolution of TensorFlow and Keras, stakeholders assumed that Keras 2.2.2 can be used with TensorFlow 1.3 without being aware of the changes in the Softmax function in Keras 2.2.2, as shown in Figure 1, which led to system failures.
Figure 1: An invalid assumption of Softmax implementation in Keras 2.2.2 and call in TensorFlow 1.3 and 1.5 Zhang2020
Many industrial cases also indicate the importance of assumptions and their management. For example, in October 2018, Lion Air Flight 610 (the aircraft is Boeing 737 MAX) crashed 13 minutes after take-off and killed all 189 people on board; in March 2019, Ethiopian Airlines Flight 302 (the aircraft is Boeing 737 MAX) crashed and ended another 157 lives. According to government reports, a critical reason for Boeing 737 MAX crashes is related to poorly managed assumptions US2019 committee2020 . The reports indicated that the aircraft manufacturer made invalid assumptions about critical system components. Specifically, invalid assumptions regarding the MCAS (Maneuvering Characteristics Augmentation System) are the root cause of crashes. The reports also highlighted the need to identify and re-evaluate important assumptions in the system.
Assumptions are related to many types of software artifacts, such as decisions, technical debt, and source code Yangsca2021 . For example, in TensorFlow, there is an assumption: “TODO: Looks like there is an assumption that weight has only one user. We should add a check here”, which induces a technical debt issue: a check should be added on whether the weight has a single user in the future. Existing research on assumptions and their management in software development usually uses experiments, surveys, and case studies to manually identify and extract assumptions through observation, questionnaires, interviews, focus groups, and documentation analysis Yangsms2018 . Since such approaches have high costs (e.g., time and resources), the comprehensiveness of the identified and extracted assumptions in these studies is often limited, leading to various problems in developing new theories, approaches, and methods of assumptions and their management in software development. In this work, to overcome the issues of manually identifying assumptions in DL framework development, we conducted an exploratory study to find a good algorithm and model to identify assumptions on GitHub. This provides a foundation for further research on assumptions and their management in the development of DL frameworks, for example, assumption classification, documentation, evaluation, and reasoning  Yangsms2018 .
The main contributions of this study are as follows.
(1) We designed a set of applicable criteria to identify assumptions and constructed a new and largest dataset (i.e., AssuEval yangpackage2023 ) of assumptions collected from GitHub projects.
(2) We evaluated the performance of seven traditional ML models, a popular DL model (ALBERT), and a large language model (GPT-3.5) to identify assumptions on the AssuEval dataset.
The remainder of the paper is organized as follows. Section 2 and Section 3 provide the related work and background knowledge, respectively, Section 4 presents the design of the evaluation on different classification models, Section 5 presents the results, Section 6 discusses the results, Section 7 clarifies the threats to the validity of the study, and Section 8 concludes the work with future directions.
2 Related work
In the field of assumptions and their management in software development, most assumptions are manually identified by researchers and practitioners. Landuyt and Joosen focused on the assumptions made during the application of a threat modeling framework (that is, LINDDUN), which allows the identification of privacy-related design flaws in the architecting phase Landuyt2020 . They conducted a descriptive study with 122 master students, and the students identified and extracted 845 assumptions from the models created by the students. In our previous work, we conducted an exploratory study as a first step in understanding assumptions in terms of their distribution, classification, and impacts based on code comments from nine popular DL framework projects on GitHub Yangsca2021 . The results show that (1) 3,084 assumptions are scattered across 1,775 files in the nine DL frameworks, ranging from 1,460 (TensorFlow) to 8 (Keras) assumptions. (2) There are four types of assumption validity: Valid, Invalid, Conditional, and Unknown and four types of assumptions based on their content: Configuration and Context, Design, Tensor and Variable, and Miscellaneous. (3) Both valid and invalid assumptions can have an impact within a specific scope (e.g., in a function) on the DL frameworks. Certain technical debt is induced when making assumptions, and source code is written and decisions are made based on assumptions. Xiong et al. studied the assumptions in the Hibernate Developer mailing list, including their expression, classification, trend over time, and related artifacts Xiong2018 . In their study, they identified and extracted 832 assumptions. Li et al. developed an ML approach Li2019 to identify and classify assumptions based on the dataset constructed by Xiong et al. Xiong2018 , which can read the data (i.e., sentences) from the dataset (i.e., a .csv file), preprocess the data (e.g., using NLTK and Word2Vec), train specific classifiers (e.g., Perception, Logistic Regression, and Support Vector Machine), and evaluate the trained classifiers (i.e., precision, recall, and F1 score).
There are also studies that focus on other types of artifacts in the development of DL frameworks and identify assumptions related to the artifacts. Zhang et al. Zhang2020 collected 4,960 real failures from a Microsoft DL platform. This platform includes multiple DL frameworks, such as TensorFlow, PyTorch, and Keras. The authors manually examined the failure messages from the 4,960 system failures and classified them into 20 categories. They also identified root causes and solutions in a sample of 400 failures, including system failures caused by assumptions. Islam et al. Islam2019 focused on bugs in DL frameworks collected from both Stack Overflow and GitHub, and studied Caffe, Keras, TensorFlow, Theano, and Torch to understand the types, root causes, and impacts of bugs, the bug-prone stage of the DL pipeline, and related antipatterns in the frameworks. They mentioned that some bugs are caused by invalid assumptions of the behavior and specifications for different APIs in different frameworks/libraries. For example, different frameworks may have similar functions, such as adjusting the size of an image (e.g., PIL in Keras and tensorflow.image.resize_images in TensorFlow). Users usually need to adjust the image size in multiple components of a DL system and may use different APIs provided by different frameworks interchangeably, assuming that they have the same behavior. However, when such an assumption is invalid, it can cause system failures. Zhang et al. Zhang2018 conducted an empirical study on 175 TensorFlow bugs collected from Stack Overflow and GitHub. They studied the root causes and symptoms of bugs and solutions for bug detection and localization. The authors also pointed out that certain system failures are caused by invalid assumptions. For example, they found that stakeholders made an invalid assumption regarding computational models, which caused the construction of an invalid TensorFlow computational graph. Specifically, a TensorFlow user used tf.assign to calculate a Fibonacci sequence and defined three nodes as as0, as1, and as2, as shown in Figure 2(a). The user assumed that the nodes followed the conventional control-flow semantics and would be computed sequentially, which is an invalid assumption, since TensorFlow does not impose a sequential order on the computation of the nodes, i.e., the nodes can be computed in an arbitrary order, as shown in Figure 2(b).
((a)) Implementation of the Fibonacci sequence calculation
((b)) Arbitrary results of the Fibonacci sequence calculation
Figure 2: An example of TensorFlow misuse caused by an invalid assumption Zhang2018
As described in the related studies mentioned above, assumptions and their management are important in the development of DL frameworks. However, existing approaches are either time-consuming or not specifically designed to identify assumptions in the development of DL frameworks.
3 Background
3.1 Machine learning algorithms
3.1.1 Perceptron
The Perceptron (Pct) algorithm is based on a threshold function as shown in Equation 1. $f(x)$ is the output, $w$ is a vector of weights, $x$ is the input, which is also a vector, and $b$ is the bias. The value of $f(x)$ is used to classify the input as type one or type two. Some studies used +1 and -1 instead of 1 and 0, which has the same rationale.
$f(x)=\begin{cases}1,ifw\cdot x+b>0\\ 0,otherwise\\ \end{cases}$ (1)
The Perceptron algorithm can also perform multiclass classification tasks, as shown in Equation 2. $x$ is the input, $y$ is the output, and $w$ is a vector of the weights. $f(x,y)$ is a feature representation function, which maps each possible input and output to a finite-dimensional feature vector. $\hat{y}$ is the resulting score, which represents the most possible of many outputs.
$\hat{y}=argmax_{y}f(x,y)\cdot w$ (2)
3.1.2 Logistic Regression
The Logistic Regression (LR) algorithm is similar to the Linear Regression algorithm, but uses a logistic function (that is, the sigmoid function) to deal with binary classification tasks, which is defined in Equation 3. $t$ is the input, and the output value is between 0 and 1 Jr2000 .
$\sigma(t)=\frac{1}{1+e^{-t}}$ (3)
Assuming that $t$ is linear, $x$ is the input vector, and $b$ is the bias (as shown in Equation 4), then the general logistic function can be described as Equation 5. $p(x)$ represents the probability that the output is in Category one and $(1-p(x))$ represents the probability that the output is in Category two.
$t=w\cdot x+b$ (4)
$p(x)=\sigma(t)=\frac{1}{1+e^{-(w\cdot x+b))}}$ (5)
LR can also be generalized to multiclass classification problems. Instead of using the Sigmoid function, LR can use the Softmax function to find the predicted probability of each category. For $i=1,....,N$, the standard Softmax function is shown in Equation 6. $N$ is the number of categories, $t$ is the input, and $t=(t_{1},t_{2},...,t_{N})\in\mathbb{R}^{N}$. The Softmax function uses an exponential function for each $t_{i}$ and ensures that the sum of the output vectors is 1.
$\sigma(t)_{i}=\frac{e^{t_{i}}}{\sum\limits_{j=1}^{N}e^{t_{j}}}$ (6)
3.1.3 Linear Discriminant Analysis
Linear Discriminant Analysis (LDA) can be used in both binary classification and multiclass classification Balakrishnama1998 . LDA maximizes the ratio of variance between classes to variance within classes in the datasets, which can ensure maximal separability Balakrishnama1998 . There are different strategies in LDA, for example: (1) maximize the ratio of variance between classes to variance within classes and (2) maximize the ratio of overall variance to variance within classes Balakrishnama1998 . Consider a binary classification problem with two datasets, the first step is to compute the mean (e.g., $\mu_{1}$ and $\mu_{2}$) of each dataset separately and the mean (e.g., $\mu_{3}$) of both datasets. Let $p_{1}$ and $p_{2}$ be the apriori probabilities of the classes, then we have Equation 7.
$\mu_{3}=p_{1}*\mu_{1}+p_{2}*\mu_{2}$ (7)
The within-class scatter is shown in Equation 8. For the binary problem, the within-class scatter is shown in Equation 9. Let $conv_{1}$ and $conv_{2}$ be the covariance of the training set and the test set. The equation of covariance is shown in Equation 10.
$S_{w}=\sum\limits_{j}p_{j}*(cov_{j})$ (8)
$S_{w}=p_{1}*cov_{1}+p_{2}*cov_{2}$ (9)
$cov_{j}=(x_{j}-\mu_{j})(x_{j}-\mu_{j})^{T}$ (10)
The between-class scatter is shown in Equation 11. Using different strategies in LDA will lead to a different calculation of the optimization criterion. After transformation of the data based on the selected strategy, we can further calculate the distance (e.g., the Euclidean distance) of the data from each class mean.
$S_{b}=\sum\limits_{j}(\mu_{j}-\mu_{3})*(\mu_{j}-\mu_{3})^{T}$ (11)
3.1.4 K-Nearest Neighbors
Fix and Hodges proposed a non-parametric method for pattern classification, which is the basis of K-Nearest Neighbors (KNN) Fix1989 . Cover and Hart expanded the KNN algorithm by adding formal properties of the rule Cover1967 . KNN can be used for both classification and regression problems. For the KNN classification, we first calculate the $k$ nearest neighbors of a data sample (e.g., the Euclidean distance). The data sample is then classified according to the vote of its $k$ nearest neighbors. Let $x_{i}=(x_{i}1,x_{i}2,…,x_{i}p)$ be a data sample. $p$ is the total number of features. The Euclidean distance between sample $x_{i}$ and $x_{j}(j=1,2,…,n)$ can be calculated using Equation 12.
$d(x_{i},x_{j})=\sqrt{(x_{i}1-x_{j}1)^{2}+(x_{i}2-x_{j}2)^{2}+...+(x_{i}p-x_{j}p)^{2}}$ (12)
3.1.5 Support Vector Machine
Support Vector Machine (SVM) is a commonly used supervised ML algorithm for data classification and regression analysis Cortes1995 . In a binary classification problem, SVM maps samples to a feature space and maximizes the width of the gap using a hyperplane between the two categories. If there is a new sample, it is mapped into the same feature space and assigned to Category 1 or 2 based on which side of the gap they fall into. SVM supports both linear classification and non-linear classification. For linear classification, taking into account $n$ points as $(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n})$, $x_{i},i=1,2,...,n$ is a p-dimensional vector, $y_{i},i=1,2,...,n$ is the label of $x_{i}$, indicating which category $x_{i}$ belongs to. A hyperplane can be described as Equation 13. The parameter $\frac{b}{||w||}$ represents the offset of the hyperplane. For non-linear classification, SVM uses kernels to map the inputs into high-dimensional spaces, including Polynominal, Gaussian radial basis function, and Sigmoid function.
$w\cdot x-b=0$ (13)
3.1.6 Naive Bayes
Naive Bayes (NB) classifiers are based on the Bayes’ theorem and assume the features are independent McCallum1998 . Let the input $x$ have n features, which is $(x_{1},x_{2},...,x_{n})$. For Category $k$, the probability is $p(C_{k}|x)$. According to Bayes’ theorem, we can obtain Equation 14. NB classifiers usually use a decision rule with the model, i.e., the most probable one is the category.
$p(C_{k}|x)=\frac{p(C_{k})p(x|C_{k})}{p(x)}$ (14)
There are different types of NB models, such as the multivariate Bernoulli model and the multinomial model, which use different strategies to calculate the probability of each category McCallum1998 .
3.1.7 Classification and Regression Trees
In a classification problem, the non-leaf nodes of a decision tree are input features, and the leaf nodes of a decision tree are categories or probability distributions over the categories. The decision tree is constructed by iteratively splitting on the basis of specific rules. Considering there are three classes and two features, Figure 3 shows an example of the structure of a decision tree. A case goes to the left or right child node only if the condition is fulfilled. The leaf nodes represent the categories.
Figure 3: Structure of a decision tree Loh2011
There are many tree learning algorithms used for classification or regression problems, such as Iterative Dichotomiser 3 (ID3) Quinlan1986 , C4.5 (which is the successor of ID3) Quinlan1993 , Classification And Regression Trees (CART) Breiman1984 , and Chi-Square Automatic Interaction Detection (CHAID) Kass1980 . These algorithms use different strategies Loh2011 . For example, CART uses the Gini index to measure how well a node separates the data in order to choose the best split of the tree Loh2011 .
3.2 BERT and ALBERT
The architecture of the BERT model is based on the original Transformer model Vaswani2017 , which is a multilayer bidirectional Transformer encoder Devlin2019 , as shown in Figure 4. The BERT model was designed for multitasks, such as natural language understanding, question-and-answer, and sentence completion. In the original paper, Vaswani et al. proposed two versions of the BERT model: BERT-Base and BERT-Large Vaswani2017 . BERT-Base has 12 transformer blocks and 12 self-attention heads, the hidden size is 768, and the total parameters are 110M; BERT-Large has 24 transformer blocks and 16 self-attention heads, the hidden size is 1024, and the total parameters are 340M Vaswani2017 . To better deal with environments with restricted computational resources, Turc et al. proposed different versions of the BERT model with smaller sizes, for example, BERT-Tiny, BERT-Mini, BERT-Small, and BERT-Medium, which reduce the number of transformer blocks and hidden size turc2019 .
Figure 4: Architecture of the Transformer model Vaswani2017
The input of the BERT model accepts a sentence or pair of sentences in a token sequence to allow the model to perform different tasks Devlin2019 . Devlin et al. used WordPiece embeddings with a vocabulary of 30,000 tokens Devlin2019 . In the embeddings, they used a specific token: [SEP] to separate the two sentences and added a learned embedding to each token to show whether a token belongs to the first sentence or the second Devlin2019 . Taking into account an input that includes two sentences: “my dog is cute” and “he likes playing”, the input embeddings are the sum of the token embeddings, the segmentation embeddings, and the position embeddings, as shown in Figure 5.
Figure 5: BERT input embeddings Devlin2019
Since the trained BERT model can successfully deal with a broad set of NLP (Natural Language Processing) tasks, many researchers tried to adapt the original BERT model, leading various models, such as XLNet YangXlnet2019 , RoBERTa LiuRoBERTa2019 , ALBERT LanALBERT2020 , and ELECTRA ClarkELECTRA2020 . These models focus on different aspects of improving the original BERT model. In this paper, we chose one of the most representative models: ALBERT, which aims to reduce parameters, lower memory consumption, and increase the training speed of the BERT model LanALBERT2020 . Compared to the BERT model, ALBERT has a similar architecture but uses factorized embedding parameterization, cross-layer parameter sharing, and inter-sentence coherence loss for optimization LanALBERT2020 . The ALBERT model has four versions: ALBERT-Base (12M parameters), ALBERT-Large (18M parameters), ALBERT-xLarge (60M parameters), and ALBERT-xxLarge (235M parameters), with much fewer parameters than the BERT model, but did a better job in certain tasks (e.g., GLUE, RACE, and SQuAD benchmarks) LanALBERT2020 .
3.3 GPT-3.5 series models
The GPT-3.5 series of models includes code-davinci-002, text-davinci-002, text-davinci-003, and gpt-3.5-turbo-0301, which can be traced back to 2020222https://platform.openai.com/docs/model-index-for-researchers. Brown et al. discovered that scaling language models could significantly enhance task-agnostic, few-shot performance, sometimes even surpassing the performance of prior state-of-the-art fine-tuning approaches Brown2020 . Then they proposed GPT-3, which comes in three versions: GPT-3 175B (175 billion parameters), GPT-3 6.7B (6.7 billion parameters), and GPT-3 1B (1 billion parameters). They have demonstrated strong performance on a variety of NLP tasks, including translation, question-answering, cloze tasks, unscrambling words, using a novel word in a sentence, and performing 3-digit arithmetic. Chen et al. proposed Codex: a GPT language model fine-tuned on publicly available code from GitHub Chen2021 . This model has only one version: Codex 12B (12 billion parameters). To evaluate its Python code-writing capabilities, they created a new dataset called HumanEval, which can be used to measure functional correctness for synthesizing programs from docstrings. In HumanEval, Codex solved 28.8% of the problems, while GPT-3 solved 0% of the problems. The authors also discovered that repeated sampling from the model is an effective way to generate working solutions to difficult prompts. With 100 samples per problem, they were able to solve 70.2% of their problems using the repeated sampling strategy. Neelakantan et al. investigated pretraining techniques Neelakantan2022 . They discovered that using contrastive pretraining on unsupervised data with a large batch size can generate high-quality vector representations of text and code. Their models achieved the best results on linear-probe classification, text search, and code search, but underperformed on sentence similarity tasks. In their work, they trained the GPT-3 unsupervised cpt-text model, which has three versions: GPT-3 unsupervised cpt-text 175B (175 billion parameters), GPT-3 unsupervised cpt-text 6B (6 billion parameters), and GPT-3 unsupervised cpt-text 1.2B (1.2 billion parameters). Stiennon et al. created a dataset of human-generated comparisons between summaries and then used it to train a model to predict the human-preferred summary Stiennon2022 . This model was also used as a reward function to fine-tune a summarization policy through reinforcement learning. This model has three versions: GPT-3 6.7B pre-train (6.7 billion parameters), GPT-3 2.7B pre-train (2.7 billion parameters), and GPT-3 1.3B pre-train (1.3 billion parameters). Their study shows that it is possible to significantly improve summary quality by training a model to optimize for human preferences. Ouyang et al. developed a model called InstructGPT, which has three versions: InstructGPT-3 175B (175 billion parameters), InstructGPT-3 6B (6 billion parameters), and InstructGPT-3 1.3B (1.3 billion parameters) Ouyang2022 . Even the InstructGPT-3 1.3B version, which has much fewer parameters than GPT-3 (175B parameters), outperformed GPT-3 in certain areas (e.g., truthfulness and reductions in toxic output generation). Furthermore, they found that fine-tuning with human feedback is a promising approach to align language models with human intent. OpenAI published a report of GPT-4 in 2023 OpenAI2023 . GPT-4 is also based on Transformer, which is the same as the GPT-3 and GPT-3.5 series of models. GPT-4 is capable of taking not only written texts, but also images. The experiments in the report demonstrate that GPT-4 is superior to existing large language models (e.g., GPT-3 and GPT-3.5 models) on a set of NLP tasks and surpasses the majority of the reported state-of-the-art systems.
As mentioned in related studies, the GPT models have evolved over time and can be used for multiple tasks, including identifying assumptions in the development of DL frameworks.
4 Study design
The objective of the study is to evaluate different classification models for the purpose of identification with respect to assumptions from the point of view of developers and users in the context of DL framework projects (i.e., issues, pull requests, and commits) on GitHub.
4.1 Research questions
RQ1: Which ML model achieves the best performance on the AssuEval dataset?
Rationale: Algorithms usually have varying performance in classification problems, leading to different trained models. For example, Naive Bayes models are a group of probabilistic classifiers that have been widely used in text categorization since 1960. The Support Vector Machine algorithm is a commonly used supervised ML algorithm for data classification and regression analysis. To answer RQ1, we chose seven popular ML algorithms to train classification models for the identification of assumptions: Perceptron (Pct), Logistic Regression (LR), Linear Discriminant Analysis (LDA), K-Nearest Neighbors (KNN), Support Vector Machine (SVM), Naive Bayes (NB), and Classification and Regression Trees (CART). These algorithms have been used in similar studies, such as the classification of non-functional requirements Abad2017 and rationale Alkadhi2018 . Li et al. also used these seven ML algorithms to identify assumptions (binary classification) in the Hibernate project Li2019 . Through this RQ, we plan to evaluate the performance of traditional ML models for identifying assumptions on the AssuEval dataset yangpackage2023 .
RQ2: What is the performance of ALBERT on the AssuEval dataset?
Rationale: In addition to ML models, DL models have also been widely used in various areas as a key component, such as image classification, question and answer, and speech recognition. For answering RQ2, we chose ALBERT to train a classification model to identify assumptions. ALBERT is widely used in many similar tasks, such as the classification of sentiments in Yelp reviews Alamoudi2021 or movie reviews Ding2021 , the classification of Japanese text Zhang2022 , and the detection of specific speech in a given sentence Vijayakumar2022 . We consider ALBERT as a representative model in DL. The results of this RQ provide a comparison of the performance in identifying assumptions between the DL model and the ML models.
RQ3: Is ChatGPT able to perform the task of identifying assumptions?
Rationale: In addition to traditional ML and DL models, large language models (LLMs) are another potential solution for identifying assumptions. In this RQ, we chose one of the most representative LLMs (i.e., ChatGPT) to explore its performance in identifying assumptions. ChatGPT is an advanced AI chatbot developed by OpenAI [^3] https://openai.com/, which can interact in a conversational way chatgptintro2022 . As reported by OpenAI, more than 100 million people used ChatGPT [^4] https://chatgptui.org/features?step=1. ChatGPT can be applied in various fields, such as answering specific questions, writing an essay, writing a song, writing code to implement a function, and code debugging chatgptintro2022 Meyer2023 . Specifically, we chose the free version [^5] https://chat.openai.com/chat of ChatGPT (based on GPT-3.5) as it is more available to the public. The results of this RQ show the performance of ChatGPT in identifying assumptions on the AssuEval dataset.
4.2 Process of study execution
We conducted the study based on a framework to identify assumptions on GitHub composed of four components: Data Collection, Dataset Management, Training, and Data Analysis, as shown in Fig. 6. Data Collection aims to (1) get information on repositories and their releases from the GitHub server and (2) collect commits, pull requests (PRs), issues, etc. from specific repositories. Dataset Management includes (1) label management (e.g., adding or editing a label), (2) labeling of raw data (e.g., commits) from specific repositories, and (3) generation of datasets into files (e.g., csv files). Training first reads and processes data from the generated datasets, constructs ML/DL models, and then trains the models. Data Analysis uses the trained models to identify assumptions on new data. We developed a tool named Assumption Miner Yangtool2023 to implement the framework and conduct the study. Details of the tool can be found in the replication package of this study yangpackage2023 .
Figure 6: Assumption identification framework for GitHub
4.3 Data collection
The process of data collection is shown in Fig. 7.
(1) Select repositories. TensorFlow is one of the most popular DL frameworks, which is widely used in many DL systems and application domains. The TensorFlow project on GitHub started in 2015, having 156,600 commits [^6] https://github.com/tensorflow/tensorflow (accessed on 2023-11-26), 23,530 PRs [^7] https://github.com/tensorflow/tensorflow/pulls?q=is%3Apr (accessed on 2023-11-26), and 38,344 issues [^8] https://github.com/tensorflow/tensorflow/issues?q=is%3Aissue (accessed on 2023-11-26) till November 2023. Therefore, in this study, we chose the TensorFlow repository on GitHub as the main source to build our dataset. Moreover, Keras is an independent repository on GitHub, which is closely related to TensorFlow. For example, in some versions of TensorFlow, Keras has been integrated inside TensorFlow as a sub-module (i.e., tensorflow.keras). Therefore, in addition to TensorFlow, we also included the Keras repository on GitHub when constructing the AssuEval dataset yangpackage2023 .
(2) Determine the scope of the data. According to similar works (e.g., Zhang2018 Ruan2019 Han2020 ) that collect and analyze the data on GitHub, in this study we focus on commits, PRs, and issues of the selected repositories.
(3) Determine the data period. Repositories on GitHub are iteratively updated. The data period to be collected is from the creation of the TensorFlow and Keras repositories till the date we started the data collection process (that is, 2023-07-19). Data after this date are not included in this study.
(4) Collect the data. We used Assumption Miner Yangtool2023 to collect the data. This tool automatically collected data from the GitHub sever using the GraphQL API [^9] https://docs.github.com/en/graphql/overview/about-the-graphql-api.
(5) Review the collected data. The first author checked the correctness and completeness of the data collected. For example, the first author visited the TensorFlow and Keras repositories on GitHub to check if the numbers of commits, PRs, and issues are consistent with the numbers in the local database.
Figure 7: Process of data collection
4.4 Dataset management
4.4.1 Construction process of the AssuEval dataset
The process of building the dataset is shown in Fig. 8, which contains 11 steps as follows.
(1) Determine the labeling scope. Not all of the information in commits, PRs, and issues is meaningful for the identification of assumptions. Therefore, after analyzing the data models provided by GitHub, the scope of the data to be collected is shown in Table 1. Details of the GraphQL API and the objects can be found on GitHub Docs [^10] https://docs.github.com/en/graphql/reference/objects.
Table 1: Scope of the GitHub data to be collected
Object Field Description
Commit message The Git commit message.
PR title The title of the PR.
PR body The body of the PR.
PR comments.body A list of comments associated with the PR.
Issue title The title of the issue.
Issue body The body of the issue.
Issue comments.body A list of comments associated with the Issue.
(2) Create labels. We created three labels, i.e., NA with a value 0, PA with a value 1, and SCA with a value 2, which represents non-assumptions, potential assumptions, and self-claimed assumptions, respectively.
(3) Conduct a pilot labeling. We conducted a pilot labeling on the TensorFlow and Keras repositories. The criteria were refined iteratively during the pilot study. For example, regarding an assumption that is a condition (e.g., “This obviously works if you can assume 0.0 is not real value in your data”), we decided not to label such data as SCA, because the assumption does not exist. Another example is that if a sentence is a question (e.g., “@yaroslavvb I’m assuming that if I run the probe op in a session together with computation of a model this would return me the peek memory usage, is that correct?”), which is about the validity of an assumption rather than whether to make an assumption, we decided to label such data as SCA. More details of the included or excluded examples can be found in Table 2, Table 3, Table 4, and Table 5.
Figure 8: Construction process of the AssuEval dataset
(4) Design criteria for labeling. After the pilot study, we formed the final criteria for labeling the data as mentioned in Section 4.4.2.
(5) Label the data from the selected repositories. The first author labeled the data following the criteria using Assumption Miner. NAs are labeled as “0”, PAs are labeled as “1”, and SCAs are labeled as “2”.
(6) Error check and discussion. The first author checked the labeled data and then asked the other authors to review the results. All authors reached a consensus on the results.
(7) Export the labeled data. This step was also done by using Assumption Miner.
(8) Remove duplicates. Since commits, PRs, and issues may share some information, and the TensorFlow repository and the Keras repository are partly overlapped (e.g., TensorFlow uses Keras functions as high-level APIs), there could be duplicates in the labeled data. In this step, we removed all duplicates of the labeled data. Note that duplicates in this study mean that sentences are exactly the same, not only the meaning of sentences, but also the words, symbols, and spaces used, which can increase the diversity of the data. For example, we considered “# 32 floats > compression factor 24.5, assuming the input is 784 floats” and “# 32 floats > compression of factor 24.5, assuming the input is 784 floats” as two different data, though they have the same meaning.
(9) Select SCAs, PAs, and NAs. According to our previous studies (e.g., Yangsca2021 and Yangtool2023 ), there are more than 1 million sentences in the TensorFlow and Keras repositories. The number of NAs is much higher than that of PAs, and the number of PAs is much higher than that of SCAs. Therefore, in this study, we selected all manually identified SCAs from the TensorFlow and Keras repositories. Then we randomly selected PAs and NAs with the same number of SCAs to construct the AssuEval dataset.
(10) Final review of the selected data. All the authors conducted a final review on the selected data and reached a consensus on the labeling results.
(11) Create the final dataset file. After review, we created the final SCA, PA, and NA dataset file, which is available at yangpackage2023 . In the dataset file, the label “0” represents NA, “1” represents PA, and “2” represents SCA.
4.4.2 Criteria of data labeling
Motivation of defining criteria for assumption identification. Although stakeholders constantly make assumptions in their work Yangsms2018 Yangsurvey2016 , the definition or characteristics (e.g., subjective) of an assumption are abstract (e.g., something is uncertain but accepted as true without evidence) Yangsms2018 Yangtool2021 . As an example, Roeller et al. claimed that “from one perspective or stakeholder, we may denote something as an assumption, while the same thing may be seen as a design decision from another perspective” Roeller2006 . This leads to difficulties in designing criteria to identify whether a sentence is an assumption in projects. In traditional assumption identification, most related studies chose to ask directly stakeholders if something is an assumption through experiments, surveys, and case studies. For example, as mentioned in the related work, Landuyt and Joosen collected 845 assumptions with 122 master students using a descriptive study Landuyt2020 . Such approaches require high costs and the number of assumptions identified is usually limited  Yangsms2018 . Since we need a large dataset, these approaches are not suitable for this study.
The SCA concept and the PA concept. As the scope of the study is to analyze the DL framework projects on GitHub, before designing the dataset construction criteria, we first classified assumptions into two types: self-claimed assumptions (SCAs) and potential assumptions (PAs). The rationale of such a classification is that a sentence cannot be identified as an assumption unless it is explicitly mentioned by stakeholders, which are SCAs. There are also some plausible cases that may or may not be assumptions, which we have to ask stakeholders to confirm, which are PAs. Taking into account a sentence in a commit, PR, or issue of a GitHub repository, an SCA is that the sentence includes an assumption and the assumption is explicitly claimed using at least one of the assumption-related terms (i.e., “assumption”, “assumptions”, “assume”, “assumes”, “assumed”, “assuming”, “assumable”, and “assumably”). For example, in this sentence “[tf/xla] fixup numbering of xla parameters used for aliasing previously, the xla argument parameter was incorrectly assumed to correspond to the index in the vector of ‘xlacompiler::argument’”, it includes an SCA of assuming that the xla argument parameter corresponds to the index in the vector of xlacompiler::argument and the developer claimed that it was an invalid assumption. A PA is that the sentence may include an assumption, which is a potential assumption that needs further confirmation from human experts. This definition covers various aspects, such as expectations, future events, possibilities, guesses, opinions, feelings, and suspicions, which can indicate PAs. We provide three examples to further explain the concept of PA. For example, in the sentence “I think the right way to create demo tensorboard instances is to simply run a tensorboard in the cloud, rather than keep maintaining this mocked-out backend.”, it includes a PA regarding thoughts on the right way to create demo tensorboard instances. In another example “The system will not crash under heavy load”, the sentence describes the future state of the system, which is uncertain and includes a PA. The third example “both false and true outputs should be considered independently” does not include assumption-related terms, but the sentence describes an expectation (i.e., “something should be”), which is a PA. After further confirmation by human experts, this PA can be transformed into an SCA or other types of software artifacts. Besides SCAs and PAs (which belong to explicit assumptions), there are also many implicit assumptions in projects (e.g., in stakeholders’ heads or requiring reasoning). Identifying implicit assumptions is much more tricky than identifying SCAs and PAs, since there are no explicit clues in the sentences and stakeholders need to infer the sentences based on the context, which involves assumption reasoning. Identifying implicit assumptions is beyond the scope of this study and we treat it as future work. Therefore, for a sentence that is neither an SCA nor a PA, we labeled it as NA.
Inclusion and exclusion criteria for SCAs and PAs. Based on the definition of the SCA concept and the PA concept, we designed the inclusion and exclusion criteria for SCAs with examples as shown in Table 2 and Table 3, and the inclusion and exclusion criteria for PAs with examples as shown in Table 4 and Table 5. Since the sentences in the commits, PRs, and issues are written by stakeholders of OSS projects, the sentences may not be standard English. For example, considering the sentence: “Perhaps @farizrahman4u’s implementation could be used instead?”, though it has a question mark, it is not the standard form of a question, and therefore, we considered such sentences as not a question. For conditions, a sentence usually has two parts: a condition and a statement that is constrained by the condition. Considering that a sentence has an SCA, if the SCA is in the condition, we labeled it as NA, while if the SCA is in the statement that is constrained by the condition, we labeled it as SCA. For alternatives, if making an SCA is only an option, we labeled it as NA (see the examples in Table 3). This may be mixed with other types of alternatives. For example, we treat “Not sure if it is necessary, but a more complicated alternative would be to assume that int size should mirror floatX.” as an SCA, since the alternative is regarding a solution to a problem, rather than whether to make an assumption. These rules also apply to the inclusion and exclusion criteria for PAs. In this study, we considered assumption identification as a multiclass classification problem, and therefore, we provided each sentence with only one label, that is, the labels are mutually exclusive. However, a sentence may be long and includes both SCAs and PAs (e.g., part of the sentence claims an SCA while the other part of the sentence describes a PA), which leads to the multiclass classification problem being a multi-label classification problem. We consider this as future work.
Table 2: Inclusion criteria for self-claimed assumptions (SCAs)
Inclusion criteria for SCAs Examples
SCA_IC1: A sentence (a) contains valid or invalid assumptions using assumption-related keywords (including “assumption”, “assumptions”, “assume”, “assumes”, “assumed”, “assuming”, “assumable”, and “assumably”) and (b) the assumptions have been made, i.e., not questions, conditions, and alternatives.
E1: For Ubuntu 21.04 the manual installation instructions are a thing of the past assuming that you have installed your NVIDIA drivers from the Canonical default apt repo.
E2: The main assumption behind the Glorot initialization is that the variance of the gradients should be the same in each layer.
E3: **Behavior**: When ‘tf.image.flip_left_right’ is applied to this ‘tf.Tensor’, the function incorrectly assumes a rank-3 shape and flips the image along the wrong axis.
SCA_IC2: A sentence (a) indicates that there are valid or invalid assumptions (without details) somewhere using assumption-related keywords and (b) the assumptions have been made, i.e., not questions, conditions, and alternatives.
E1: # Check input assumptions set after layer building, e.g. input shape.
E2: We tried to add lazy caching of the alternative row-partitioning tensors, but had to roll it back because it breaks some assumptions made by tf.cond.
E3: [XLA] Fix invalid assumption in HloComputation::CloneWithReplacements.
SCA_IC3: A sentence (a) contains assumptions or indicates that there are assumptions (without details) somewhere using assumption-related keywords, (b) the assumptions are not conditions and alternatives, but they are in a question, and (c) the question does not concern whether something is an assumption or making certain assumptions, that is, such assumptions exist.
E1: Assuming that it is related, is there any caching involved when reading the .tfrecord?
E2: Can you update us on when all these features will be available and in which tensorflow version (only 2.0 I assume)?
E3: @yaroslavvb I’m assuming that if I run the probe op in a session together with computation of a model this would return me the peek memory usage, is that correct?
SCA_IC4: A sentence refers to warnings that you should not make certain assumptions using assumption-related keywords.
E1: This CL generalizes the expansion such that it does not assume that ‘prefix’ is a constant - it works for both constants and non-constants.
E2: Do not assume Node.in_edges() is sorted by dst_input.
E3: Update the parser test to check for unique’d hoisted map to be present but without assuming any particular order.
Table 3: Exclusion criteria for self-claimed assumptions (SCAs)
Exclusion criteria for SCAs Examples
SCA_EC1: A sentence does not include any assumption-related keywords.
E1: Rename the pass to merge assuming ops.
E2: The improved version converts to a dynamic compatible form while restricting to static operations as often as possible.
E3: The problem occurs because PyYAML can’t recognize numpy’s data types.
SCA_EC2: A sentence includes assumption-related keywords, and the keywords are used as the name of folders, files, classes, functions, variables, etc.
E1: old ātomicCAS(address_as_ull, assumed, 0059 _double_as_longlong(val + 0060 longlong_as_double(assumed)))
E2: ga_uint old, assumed, sum, new
E3: Update ’( assuming’ to ’(assuming’.
SCA_EC3: A sentence includes assumptions-related keywords, and assumptions are conditions or alternatives.
E1: If they are independent, or you make the assumption that they are, then you can do ‘-log(P(output1=1|data) * P(output2=c|data))’ which can further decompose into ‘-log(P(output1=1|data))-log(P(output2=c|data))’.
E2: This obviously works if you can assume 0.0 is not real value in your data.
E3: So, if you assume independence between the two classifications, the linear combination of the two losses is indeed “minimizing joint negative log likelihood”.
SCA_EC4: A sentence includes assumption-related keywords, and the sentence is only about asking whether something is an assumption or whether to make certain assumptions, i.e., such assumptions do not exist.
E1: Are there any assumptions made by this method regarding the data?
E2: Can I assume that the parameter ‘y_pred’ in ‘custom_objective(y_true, y_pred)’ is just the list generated by my ‘generate_batch_data function’ and I can treat it as a normal list?
E3: What is the assumed shape of ’weights’?
Table 4: Inclusion criteria for potential assumptions (PAs)
Inclusion criteria for PAs Examples
PA_IC1: A sentence (a) is not an SCA, (b) contains valid or invalid expectations, future events, possibilities, guesses, opinions, feelings, or suspicions, and (c) the expectations, future events, etc. are not questions, conditions, and alternatives.
E1: If something/somewhere assumes that exceptions are available, things may not work as expected.
E2: Theano tile() expects Python int, so casting from numpy.int32 to Python int. (#4330)
E3: My original docs only included the dimensions of the parameters (no batch dim) and were correct, but I think its better to change the functions to reflect the current docs.
PA_IC2: A sentence (a) is not an SCA, (b) contains valid or invalid expectations, future events, possibilities, guesses, opinions, feelings, or suspicions, (c) the expectations, future events, etc. are not conditions and alternatives, but are in a question, and (d) the question is not regarding whether to make expectations, guesses, opinions, feelings, or suspicions, whether future events are going to happen, or whether something is possible.
E1: Black image, seems a big wast to me, can I input another shape tensor instead?
E2: It would force people to clean up their existing code, and their code would get better. (Also, are there any concrete examples of existing code that assumes the current broken behavior?)
E3: an unpinned copy holds that mutex for a long time (until the copy completes?), potentially blocking other, unrelated work on the gpu.
Table 5: Exclusion criteria for potential assumptions (PAs)
Exclusion criteria for PAs Examples
PA_EC1: A sentence is an SCA.
See the examples in Table 2.
PA_EC2: A sentence (a) includes expectations, future events, possibilities, guesses, opinions, feelings, or suspicions, while (b) they are conditions or alternatives.
E1: Allows fit() when call’s signature looks something like call(x, training=true).
E2: Raise exception if unexpected keys are found in the padding dict.
E3: “It is diffcult or impossible to do obsure(very unique) things in Keras”, this is what an expert in out department has suggested me.
PA_EC3: A sentence (a) includes expectations, future events, possibilities, guesses, opinions, feelings, or suspicions, while (b) the sentence consists of only asking whether to make expectations, guesses, opinions, feelings, or suspicions, whether future events are going to happen, or whether something is possible.
E1: hi @abattery, would you mind taking a look at this?
E2: it’s not legal for collective-permtue to write to the same replica twice, because then which writer is supposed to win?
E3: make the community-builds link more prominent @andrew-leaver what do you think of this?
PA_EC4: A sentence does not include any expectations, future events, possibilities, guesses, opinions, feelings, or suspicions.
E1: Added support for padded_batch and fixed comments.
E2: Fix TypeError positional argument when used conjointly with tf-addons wrappers.
E3: * When use tensorflow as backend, let batch norm run into fused batch norm as much as possible, which has better performance.
4.5 Training
We deployed Assumption Miner locally to train the selected ML models and ALBERT. The configuration of the computer we used for the training is: (1) CPU: 12th Gen Intel(R) Core(TM) i7-12700F, (2) Memory: 32GB, (3) Hard Disk: 500GB SSD, (4) GPU: NVIDIA GeForce RTX 4060 Ti.
4.5.1 Data processing
We first split the dataset into a training set (80%) and a testing set (20%), which is a common strategy used in many classification-related tasks (e.g., Jasim2021 Helber2019 Durden2021 ). When loading the data, we transformed all the loaded data into the lower case. Then we used the same vocabulary with a size of 30,000 as in ALBERT LanALBERT2020 , and we used SentencePiece to tokenize the data in the training set and the test set Kudo2018 . The vocabulary (“30k-clean.vocab” and “30k-clean.model”) can be found in the replication package of this study yangpackage2023 .
4.5.2 Model construction and training
For answering RQ1, we used scikit-learn to implement and train the seven ML models. The configurations of each model are shown in Table 6.
Table 6: Configurations of training the ML models
Pct LR SVM CVAT
tol 1e-3 tol 1e-4 tol 1e-3 criterion gini
penalty None penalty l2 coef0 0.0 splitter best
l1_ratio 0.15 l1_ratio None cache_size 200 max_depth None
fit_intercept True fit_intercept True break_ties False min_samples_split 2
max_iter 1,000 max_iter 10,000 max_iter -1 min_samples_leaf 1
verbose 0 verbose 0 verbose False min_weight_fraction_leaf 0
validation_fraction 0.1 solver lbfgs kernel rbf max_features None
n_jobs None n_jobs None decision_function_shape ovr max_leaf_nodes None
random_state 0 random_state None random_state None random_state None
alpha 0.0001 dual False gamma scale min_impurity_decrease 0
class_weight None class_weight None class_weight None class_weight None
eta0 1.0 intercept_scaling 1 degree 3 ccp_alpha 0.0
early_stopping False multi_class auto probability False min_impurity_split None
n_iter_no_change 5 C 1.0 C 1.0
warm_start False warm_start False shrinking True
shuffle True
LDA KNN NB
tol 1e-4 n_neighbors 5 alpha 1.0
shrinkage None weights uniform binarize .0
priors None algorithm auto fit_prior True
n_components None leaf_size 30 class_prior None
store_covariance False p 2
covariance_estimator None metric minkowski
solver svd metric_params None
n_jobs None
For answering RQ2, we used the TensorFlow framework to train the ALBERT model. The configurations of the ALBERT model are shown in Table 7. The optimizer is Adam, the training batch size is 64, and the maximum sequence length of each sentence is 128. We trained the model to 10,000 epochs with an initial learning rate of 2e-5.
Table 7: Configurations of training the ALBERT model
Item Value
attention_probs_dropout_prob 0
hidden_act gelu
hidden_dropout_prob 0
embedding_size 128
hidden_size 768
initializer_range 0.02
intermediate_size 3,072
max_position_embeddings 512
num_attention_heads 12
num_hidden_layers 12
num_hidden_groups 1
net_structure_type 0
gap_size 0
num_memory_blocks 0
inner_group_num 1
down_scale_factor 1
type_vocab_size 2
vocab_size 30,000
To answer RQ3, we used the pre-trained ChatGPT model without retraining or further fine-tuning. OpenAI trained ChatGPT using Reinforcement Learning from Human Feedback (RLHF) chatgptintro2022 . The ChatGPT training process is shown in Figure 9 chatgptintro2022 . The original version of ChatGPT is fine-tuned from a model in the GPT-3.5 series, which is free and completed training in early 2022 chatgptintro2022 .
We provided ChatGPT with the definitions of SCA, PA, and NA and the criteria to identify SCAs, PAs, and NAs mentioned in Section 4.4.2. Then we provided each sentence of the test set to ChatGPT and asked ChatGPT to classify it as SCA, PA, or NA.
Figure 9: Training process of ChatGPT chatgptintro2022
4.6 Data analysis
Since the AssuEval dataset (see Section 4.4) has three labels, the confusion matrix of the dataset is shown in Figure 10. $n_{i}j,i=1,2,3,j=1,2,3$ represents the number of results with respect to different conditions.
Figure 10: Confusion matrix of the AssuEval dataset
True positive (TP) is a number of positive labels that are correctly identified as positive. True negative (TN) is a number of negative labels that are correctly identified as negative. False positive (FP) is a number of negative labels that are incorrectly identified as positive. False negative (FN) is a number of positive labels that are incorrectly identified as negative. The TP, TN, FP, and FN of each label is shown in Figure 11.
Figure 11: TP, TN, FP, and FN of each label
We used accuracy, precision, recall, and f1-score to measure the classification performance of different models, as shown in Equation 15.
$\begin{split}&Accuracy=\frac{TP+TN}{TP+FP+FN+TN}\\ &Precision=\frac{TP}{TP+FP}\\ &Recall=\frac{TP}{TP+FN}\\ &F1-score=\frac{2*Precision*Recall}{Precision+Recall}\end{split}$ (15)
Since the problem is multiclass classification and the dataset is balanced, we used a macro average strategy for precision and recall calculations Grandini2020 . Taking into account that the precision of SCAs, PAs, and NAs is $P_{SCA}$, $P_{PA}$, and $P_{NA}$ and that the recall of SCAs, PAs, and NAs is $R_{SCA}$, $R_{PA}$, and $R_{NA}$, then the precision and recall are shown in Equation 16.
$\begin{split}&Precision_{macro}=\frac{P_{SCA}+P_{PA}+P_{NA}}{3}\\ &Recall_{macro}=\frac{R_{SCA}+R_{PA}+R_{NA}}{3}\\ \end{split}$ (16)
5 Results
5.1 Demographics of the dataset
As a result, we collected 150,877 commits, 23,169 PRs, and 37,560 issues from the TensorFlow repository and 8,313 commits, 6,491 PRs, and 11,764 issues from the Keras repository. On the basis of the collected data, we further constructed the AssuEval dataset containing 5,118 SCAs (1,186 from Keras and 3,932 from TensorFlow), 5,118 PAs (from Keras), and 5,118 NAs (4,998 from Keras and 120 from TensorFlow) as shown in Figure 12. As mentioned in Section 4.4.1, the number of NAs is much higher than that of PAs, and the number of PAs is much higher than that of SCAs. Therefore, in this study, we selected all manually identified SCAs (that is, 5,118) from the TensorFlow and Keras repositories. Then we randomly selected PAs and NAs with the same number of SCAs to construct the AssuEval dataset.
Figure 12: Numbers of SCAs, PAs, and NAs in the AssuEval dataset selected from the TensorFlow and Keras repositories
The sources of SCAs, PAs, and NAs include commits (598 SCAs, 542 PAs, and 4171 NAs), PRs (883 SCAs, 4044 PAs, and 675 NAs), and issues (3637 SCAs, 532 PAs, and 272 NAs), as shown in Figure 13. Specifically, we collected data from (1) message of commits (598 SCAs, 542 PAs, and 4171 NAs), (2) title (6 SCAs), body (200 SCAs, 1173 PAs, and 197 NAs), and comments.body (677 SCAs, 2871 PAs, and 478 NAs) of PRs, and (3) title (15 SCAs, 532 PAs, and 163 NAs), body (978 SCAs and 45 NAs), and comments.body of issues (2644 SCAs and 64 NAs), as shown in Figure 14.
Figure 13: Numbers of SCAs, PAs, and NAs in the AssuEval dataset across commits, PRs, and issues
Figure 14: Numbers of SCAs, PAs, and NAs in the AssuEval dataset across message of commits and title, body, and comments.body of PRs and issues
We also counted the number of words for each data item (punctuation masks in a sentence were not counted). The numbers of words for each SCA, PA, and NA range from 2 to 276, 2 to 102, and 1 to 180, respectively. To keep the figure short, we arrange the numbers into 11 intervals, as shown in Figure 15. The length of the first ten intervals is 10 (e.g., 1-10 means the sum of data items with a number of words between 1 and 10). In the figure, we can see that most of the SCAs (4,956 out of 5,118, 96.83%), PAs (5,051 out of 5,118, 98.69%), and NAs (5,108 out of 5,118, 99.80%) have less than 50 words.
Figure 15: Summary of the numbers of words in each data item in the AssuEval dataset
5.2 Results of RQ1
The confusion matrices of the Pct, LR, LDA, KNN, SVM, NB, and CART models are shown in Figure 16. The accuracy, precision, recall, and f1-score of each model are shown in Table 8. The CART model achieves the best accuracy, precision, recall, and f1-score for identifying SCAs, PAs, and NAs, while the worst performance of identifying SCAs, PAs, and NAs is the Pct model.
Figure 16: Confusion matrices of the ML models
Table 8: Performance of the seven ML models of assumption identification on the AssuEval dataset
Model Accuracy Precision Recall F1-score
Pct 0.4122 0.4205 0.4110 0.4103
LR 0.5226 0.5235 0.5245 0.5191
LDA 0.5106 0.5176 0.5124 0.5086
KNN 0.4676 0.4767 0.4703 0.4577
SVM 0.5448 0.5454 0.5454 0.5453
NB 0.4881 0.4796 0.4889 0.4674
[…]
5.3 Results of RQ2
The confusion matrix of the ALBERT model is shown in Figure 17. The accuracy, precision, recall, and f1-score of the ALBERT model are 0.9590, 0.9585, 0.9585, and 0.9584, respectively. Compared to the best model (i.e., the CART model) in RQ1, although training the ALBERT model requires much more resources (e.g., seconds vs. hours for training), the ALBERT model performs much better on the same dataset.
Figure 17: Confusion matrices of the ALBERT model
5.4 Results of RQ3
The confusion matrix of ChatGPT is shown in Figure 18. The accuracy, precision, recall, and f1-score of ChatGPT are 0.6223, 0.6476, 0.6240, and 0.6211, respectively, which is better than the CART model but much worse than the ALBERT model. Specifically, if we consider whether ChatGPT can be used to identify assumptions but not further distinguish between SCAs and PAs, that is, if we simply consider both SCAs and PAs as assumptions and therefore treat the problem as a binary classification problem (i.e., assumptions and non-assumptions), then ChatGPT performs better compared to the multiclass classification problem, with accuracy: 0.7877, precision: 0.7643, recall: 0.7880, and f1-score: 0.7711.
Figure 18: Confusion matrices of ChatGPT
6 Discussion
6.1 Performance of the models
ML models, DL models, and LLMs have different strategies to identify assumptions. Since DL models have been shown to be more efficient than ML models in many classification tasks, it is not surprising that the ALBERT model performs best in this task. However, as the most popular LLM, the performance of ChatGPT is much worse than the performance of the ALBERT model (f1-score: 0.6211 vs. 0.9584). One reason is that the version of ChatGPT we used in this study is based on the GPT-3.5 model, which was trained by OpenAI for general purposes, while the ALBERT model was trained based on the AssuEval dataset. Fine-tuning ChatGPT specifically for assumption identification could improve the performance. Moreover, although the ALBERT model achieves the highest performance, it also has limitations, such as the length of the input. In this study, if the length of a sentence is more than 128 words, it will be truncated when using the ALBERT model to identify assumptions. Therefore, the ALBERT model trained in this study can only be used to identify assumptions at the sentence level. Compared to the ALBERT model, the length limit (4096 tokens, including both input and output) of ChatGPT is much higher. In the context of ChatGPT, a token is a chunk of text that can be as short as one character or as long as one word, depending on the language and context. Tokens include words, punctuation marks, and other elements. For example, if the input sentence is “How are you?”, then ChatGPT counts it as [“How”, “are”, “you”, “?”], which has four tokens. Therefore, ChatGPT can be used to identify assumptions at the sentence level, paragraph level, or even document level.
Since ChatGPT is not only a classification model, we also asked the rationale when ChatGPT was identifying assumptions from the sentences. After analyzing the data, we can see that ChatGPT has made many errors in the classification task. Ten examples of incorrect classification with the rationale by ChatGPT are shown in Table 9. For E1, E2, E5, E7, and E9, the raw data qualify the SCA_IC1 criterion, which should be an SCA. In E1, although ChatGPT identified the keyword “assuming”, it misunderstood the sentence as a question. In E2, ChatGPT considered the sentence as “addressing the reader”, which is not true. In E5, the sentence includes “assume”, which is an assumption-related keyword, while ChatGPT said that the sentence does not use assumption-related keywords. In E7, ChatGPT explained the assumption as speculation and considered that it has to be confirmed. In E9, ChatGPT only saw part of the sentence: “something causes problems”. However, the sentence explicitly claimed an assumption that TensorFlow Probability’s JAX backend assumes properties like ‘numpy.int32 == tf.int32’ hold. For E3, ChatGPT identified an assumption-related keyword: “assumes” in the sentence, which does not exist. For E4 and E6, the sentence qualified the PA_IC1 criterion, while ChatGPT did not identify anything in the sentence. For E8 and E10, they are statements of facts, while ChatGPT considered them as assumptions. The reason is that sometimes ChatGPT would imagine the context of the sentence, even though we clearly asked it not to do that.
Table 9: Rational of assumption identification by ChatGPT
No. Raw data Ground truth ChatGPT Rationale by ChatGPT
E1
I am assuming there’s not a memory problem since on cpu it runs fine and starts training rather quickly, so I can see the progress plus when running on cpu the allocated RAM doesn’t go beyond 3GB.
SCA
PA
The sentence includes the term “assuming,” but it involves a question about the existence of a memory problem. Since the question is not explicitly an assumption, it falls into the category of Potential Assumption (PA).
E2
@petewarden Assume you’re new to TensorFlow.
SCA
NA
The sentence contains the term “Assume,” but in this context, it appears to be addressing the reader rather than making an assumption. Therefore, it is classified as Not an Assumption (NA).
E3
Remove “at least 2D” rank expansion in fit/predict/evaluate.
NA
SCA
The sentence mentions “assumes” and describes a specific assumption about the rank expansion. It meets the inclusion criteria for SCAs.
E4
…I think it would be just silly to merge the base class ‘RecurrentAttentionCellWrapperABC’ with the specific implementation ‘MixtureOfGaussian1DAttention’ as it describes and implements a highly relevant and general abstraction level for attention mechanisms.
PA
NA
The sentence discusses the merging of code without explicitly containing assumptions or potential assumptions.
E5
I assume it means the code that Theano is generating has if statements nested too deeply.
SCA
PA
The sentence indicates the author’s assumption or interpretation (“I assume it means the code that Theano is generating has if statements nested too deeply”). It does not use assumption-related keywords but expresses a potential assumption.
E6
Now it seems to be a pain point to repeat these image_dim_ordering() tests.
PA
NA
The sentence discusses a pain point related to testing but does not convey any assumptions or expectations. It is classified as Not an Assumption (NA).
E7
This sort of question is better asked on stackoverflow, assuming that it’s user error and not a completely broken model.
SCA
PA
The sentence involves a question and speculation (“assuming that it’s user error”), indicating a potential assumption that needs further confirmation.
E8
I have tried to keep the output format for the new function as close to the existing ‘categorical_accuracy()’ function as possible.
NA
SCA
The sentence explicitly mentions keeping the output format close to an existing function, indicating an assumption about maintaining similarity with the categorical_accuracy() function.
E9
This largely works fine, but causes problems for TensorFlow Probability’s JAX backend, which assumes properties like ‘numpy.int32 == tf.int32’ hold.
SCA
PA
The sentence describes a situation (“causes problems”) without explicitly stating assumptions. It suggests a potential issue without confirming assumptions.
E10
- generalized for any lahead/tsteps combination
NA
PA
The sentence is a comment related to generalization and does not explicitly state assumptions. It indicates a consideration without confirming assumptions.
6.2 Implications to researchers
Assumption documentation. In the data collection phase of this study, we can see that assumptions are scattered in various sources. Therefore, we recommend establishing best practices for documenting assumptions in DL framework projects. For example, if assumptions and their impacts can be explicitly stated, it would make it easier for practitioners to understand and manage these assumptions over time.
Assumption classification. As shown in the AssuEval dataset, there are various types of assumptions in DL framework development. In our earlier work, we identified 3,084 SCAs from code comments in over 50,000 files of nine popular DL frameworks (e.g., TensorFlow, Keras, and PyTorch) on GitHub Yangsca2021 , and analyzed the types of the identified SCAs. Researchers could work on developing a comprehensive taxonomy for classifying assumptions in DL framework development based on the taxonomy in our previous work with the dataset in this study.
Assumption evaluation. We also observed that there are critical assumptions that significantly impact the project and less crucial ones that might have minimal effects. Researchers need to develop the approach to assess the potential impact of different assumptions. This can help prioritize resources towards addressing the most critical assumptions that could influence decision-making, code structure, and overall project success.
Fine-tuning the pre-trained DL and large language models. In this study, ALBERT has shown the best performance in the identification of assumptions at the sentence level. Researchers can further investigate and develop fine-tuning strategies to improve the performance of models such as ALBERT and ChatGPT. Moreover, this finding might be generalized to decision identification, technical debt identification, etc. We need further exploration of transformer models in both assumption identification and other related classification tasks in DL framework development.
6.3 Implications to practitioners
Awareness of assumptions. Practitioners should be actively aware of the assumptions underlying the DL framework projects. This awareness is crucial to make informed decisions, anticipate potential challenges, and ensure the robustness of the developed solutions. For example, when users raise issues of the system, practitioners are encouraged to discuss and identify if there are assumptions behind the issues, and analyze their impacts on existing assumptions or other types of software artifacts. The results of the assumption analysis can be used for more robust decision-making and solutions of the issues.
Assumption documentation. To address the problems caused by scattered assumptions in various resources, practitioners need to emphasize clear and thorough documentation of assumptions in, for example, code, comments, issues, and project discussions. This will help the development team and future contributors understand the rationale behind certain design decisions and anticipate potential risks caused by assumptions.
Assumption evaluation. In the data collection phase of this study, we can see that new assumptions arise throughout the development life cycle. Since assumptions may impact other assumptions and other types of software artifacts, there is a need to review the existing ones to check, for example, if they are valid. Regular reviews ensure that the team remains proactive in managing assumptions. Moreover, since not all assumptions are important, practitioners need to understand which assumptions have a substantial impact on the project’s success and focus resources on managing and validating those assumptions appropriately.
Model selection. The selection of ML models, DL models, and LLMs is important, which should be based on the specific needs of the task at hand. According to the study findings, practitioners may lean towards using the ALBERT model when dealing with sentence-level tasks, such as assumption identification. However, practitioners should also be aware that although the ALBERT model is a lightweight transformer-based model, it has certain computational demands. Practitioners should weigh the benefits of performance against the complexity and resource requirements of these models, especially in practical software development. In addition, practitioners can also consider incorporating the ALBERT model into NLP pipelines for tasks that involve analyzing or categorizing individual sentences. Despite ChatGPT’s relatively lower performance (with an f1-score of 0.6211 compared to ALBERT’s 0.9584) in classifying assumptions within our study, practitioners may still find it beneficial to incorporate ChatGPT for understanding the context of assumptions (e.g., the rationale of making certain assumptions). ChatGPT’s capacity to generate responses and comprehend the context of assumptions can effectively be used in assumption extraction (e.g., extraction of the relationships between assumptions and requirements, design decisions, etc.) and assumption reasoning (e.g., reasoning of the conflicts between assumptions). This is the next step in sentence-level assumption analysis, which can complement the strengths of the ALBERT model, allowing for a more comprehensive understanding of assumptions.
7 Threats to Validity
We discuss the potential threats to the construct validity, external validity, and reliability of this study according to the guidelines proposed by Runeson and Höst Runeson2012 .
Construct validity reflects the degree to which the research questions and the operational measures studied are consistent Runeson2012 . A potential threat comes from the construction of the AssuEval dataset yangpackage2023 . For example, researchers might have a different understanding of the assumption concept, leading to different labels for the same sentence. To alleviate this threat, we conducted a pilot labeling on the TensorFlow and Keras repositories, and the criteria were iteratively refined during the pilot study. In the formal labeling process, all authors reviewed the results. To eliminate personal bias, conflicts and disagreements about labeling results were discussed and resolved among the authors. Moreover, as mentioned by OpenAI, ChatGPT is context dependent. For the same question, the answers may be different in various contexts. Therefore, there is a risk that ChatGPT may forget the definitions and criteria when identifying assumptions in sentences sequentially. To reduce this threat, the sentences provided to ChatGPT were separated in different chats with the same context.
External validity concerns the generalization of the findings Runeson2012 . In this study, we only selected two popular DL framework development projects to evaluate the performance of the selected models in identifying assumptions. However, the selected projects: TensorFlow and Keras are representative in DL framework development, and therefore, the results can be generalized to other DL framework projects. Moreover, assumptions may be scattered in different sources (e.g., commits, PRs, issues, code comments, and documentation), while in this study, we selected three of them: commits, PRs, and issues). To reduce the threat, we followed the guidelines proposed by Stiennon et al. Stiennon2022 to calculate and construct a representative dataset used to evaluate the performance of the selected models in the identification of assumptions. The results on the dataset can be generalized to the whole TensorFlow project. Finally, the selected algorithms are representative in ML, DL, and LLM. The results can be generalized to other related algorithms.
Reliability focuses on whether the study would produce the same results when other researchers replicate it Runeson2012 . To alleviate this threat, we iteratively discussed the research protocol that was confirmed by all authors, and we provide the details of the study design in Section 4. To facilitate replication of our work and mitigate the threat, we provide a replication package, which is available online yangpackage2023 . Moreover, ChatGPT is built on LLMs that use DL neural networks. How ChatGPT generates answers for a specific question is a black-box for its users. Therefore, we repeated our questions for ChatGPT and found that for the same question in the same context, although ChatGPT did not provide the exact same answer every time, the meaning of the answers was similar. This improves the reliability of the study. However, we note that ChatGPT has evolved over time and that users cannot access its historical versions. Therefore, for future versions of ChatGPT, they may have different answers for the questions, which can introduce reliability threats.
8 Conclusions
Stakeholders constantly make assumptions in the development of deep learning (DL) frameworks. These assumptions are related to various types of software artifacts (e.g., requirements, design decisions, and technical debt) and can turn out to be invalid, leading to system failures. Existing approaches and tools for assumption management usually depend on manual identification of assumptions. However, assumptions are scattered in various sources (e.g., code comments, commits, pull requests, and issues) of DL framework development, and manually identifying assumptions has high costs (e.g., time and resources).
To overcome the issues of manually identifying assumptions in DL framework development, we constructed a new and largest dataset (i.e., AssuEval) of assumptions collected from the TensorFlow and Keras repositories on GitHub; explored the performance of seven ML models (i.e., Pct, LR, LDA, KNN, SVM, NB, and CART), a popular DL model (i.e., ALBERT), and an LLM (i.e., ChatGPT) for identifying assumptions on the AssuEval dataset yangpackage2023 . The experiment results show that: ALBERT achieves the best performance (f1-score: 0.9584) of identifying assumptions on the AssuEval dataset, which is much better than the other models (the 2nd best f1-score is 0.6211, achieved by ChatGPT). Though ChatGPT is the most popular large language model, we do not recommend using it to identify assumptions in DL framework development because of its low performance on the task. Fine-tuning ChatGPT specifically for assumption identification could improve the performance.
For the next steps, we consider the following directions: (1) Types of assumptions: Implement the trained ALBERT model in the Assumption Miner tool and use it to identify and classify SCAs and PAs in DL framework projects. The classification can be based on various dimensions, such as the content and validity of assumptions, whether an assumption is violated, or whether an assumption has details. This helps us to understand the types of assumptions that exist in projects, which can be the basis for a further detailed analysis of these assumptions. (2) Context analysis: In this study, we only considered assumptions at the sentence level since the focus is only on identifying assumptions without further analysis. However, the context of assumptions may provide further information that can help stakeholders better understand, for example, why stakeholders made such assumptions, which is useful to explore the rationale for making assumptions in DL framework development. (3) Relationship analysis: Develop methods and tools to analyze the relationships between assumptions and other types of related software artifacts. This helps stakeholders to understand the importance of assumptions in projects and how assumptions impact each other and other types of software artifacts.
Acknowledgments
This work is funded by Shenzhen Polytechnic University with Grant No. 6022312043K, State Key Laboratory for Novel Software Technology at Nanjing University with Grant No. KFKT2022B37, the National Natural Science Foundation of China (NSFC) with Grant No. 62172311, and the Special Fund of Hubei Luojia Laboratory.